### v2

@@ 2 Problem description
We assume that researchers have access to estimators (and their variance) of $\theta$ denoted as $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$. We impose no restrictions on how $\tilde{\theta}^{\mathrm{obs}}$ is formed: it can based on arbitrary exclusion restrictions implied by an economic or statistical model. However, $\tilde{\theta}^{\mathrm{obs}}$, defined as observational estimates, have unknown biases collected in a vector $b\in\mathbb{R}^{p}$. Examples include an estimate from an instrumental variable regression that may fail the exclusion restriction, or an estimate from a country different from the one of interest that may lack external validity. Researchers can also collect experimental evidence for a subset of parameters, producing estimates whose bias equals zero by design.
@@
Here, $(\mathcal{E},\Sigma(\mathcal{E}))$ characterizes the experimental design (which parameters are learned and with what precision). When clear, we omit the subscript $\Sigma$ on $\tilde{\theta}^{\mathrm{exp}}$. We assume $\Sigma(\mathcal{E})$ is known, which is standard in experimental planning (Gerber and Green, 2012) (in practice, any consistent estimator, e.g., from pilot studies would suffice).
@@
We write compactly $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$ indicating $\mathcal{D}$ the set of feasible design (i.e., choice of $\mathcal{E}$ and $\Sigma(\mathcal{E})$) and estimators (i.e., $\gamma$). The set $\mathcal{D}$ can be arbitrary as long as (i) each feasible $\Sigma$ has uniformly bounded entries and is strictly positive definite; (ii) $\gamma_{j}=1$ for all $j\in\mathcal{E}$ is one admissible choice (among potentially other values).
We let $\mathcal{E},\Sigma(\mathcal{E})$ be within arbitrary constraint sets that may encode feasibility or budget constraints. In our framework, restrictions on the experiments researchers can run $\mathcal{E}$ and on their precision $\Sigma(\mathcal{E})$ can take any desired form. In most applications, we think of such constraints as arising from fixed costs of running an experiment and/or constraints on their power (Duflo et al., 2007; Athey and Imbens, 2017; List et al., 2011). We assume that $\Sigma(\mathcal{E})$ is uniformly bounded, which implies that once we commit to learn a set of experimental parameters $\tilde{\theta}_{\mathcal{E}}^{\mathrm{exp}}$, their variance is finite, and strictly positive definite, therefore assuming that the variance is bounded away from zero. [^5] For this latter condition to hold in an asymptotic framework with growing sample size, $\theta$ and $\tilde{\theta}$ can be defined without loss as parameters of interest after appropriately rescaling by the square-root of the sample size; in this case $\Sigma$ denotes the asymptotic variance. See Section 5.1 for details. The shrinkage estimator can be unrestricted (i.e., $\gamma$ can take any value on the real line) or restricted (when for example researchers aim to only use experimental evidence when available, setting $\gamma_{j}=1$ for $j\in\mathcal{E}$).
We extend our framework to more general methods of moments estimators in Section 4.
Remark 1 (Structure of $\Sigma$).
Because observational and experimental moments are typically computed from disjoint samples, in practice, it is natural to model the covariance as block diagonal, $\Sigma=\operatorname{diag}(\Sigma_{\text{obs}},\Sigma_{\text{exp}})$. The first block corresponds to observational estimates (fixed and not subject to design choices), while the second corresponds to experimental estimates, and it is a function of the sample size allocated to treatment arms. This is a special case of our framework which can be encoded in the constraint set $\mathcal{D}$. ∎
Research questions
@@
Choosing $\gamma$ (step 1) for a fixed design follows a long-standing tradition in econometrics and statistics (Armstrong et al., 2024; Andrews and Shapiro, 2021; Donoho, 1994; Athey et al., 2025). We study this question here in combination with the experimental design problem, step 2 and 3; this, as we show, will change the underlying optimization problem (also for $\gamma$). The central challenge is that performance depends on the unknown bias $b$.
Mapping $\tau$
A key feature of our framework is that researchers explicitly parameterize $\theta$ and $\tau(\cdot)$, thereby pre-specifying and making transparent which biases drive the design problem; we see this as an advantage because it clarifies the role of the experiment in complementing existing evidence. The sensitivity of $\tau$ to $\theta$, captured by its gradient $\omega(\theta)=\partial\tau(\theta)/\partial\theta$, determines how the (first-order) bias propagates onto the final estimand. For simplicity, we impose exact linearity, $\tau(\theta)=\omega^{\top}\theta$; all results extend to smooth, nonlinear $\tau$ via a first-order expansion, as discussed below.
@@
Assumption 2 holds exactly when $\tau(\cdot)$ is linear and serves as a first-order approximation for smooth non-linear $\tau(\cdot)$. Section 5.1 (and Example 3) extend the framework to non-linear estimands encompassing as special cases the local-misspecification setups in Andrews et al. (2020); Armstrong and Kolesár (2021). For non-linear $\tau$, we replace $\omega$ with the gradient of $\tau$ evaluated at a preliminary observational estimate, i.e., $\omega\;=\;\frac{\partial\tau(\theta)}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}\;+\;o_{p}(1),$ which is typically available before the experiment is conducted. That is, building on references above, $\omega$ captures the first-order effect of the observational bias $b$ on the bias of $\hat{\tau}$. [^6] To see how this assumption holds under a first-order approximation, define $\theta$ the parameter rescaled by $\sqrt{n}$ (and similarly its bias $b_{n}$, indexed by $n$), where $n^{-1/2}$ is the rate of convergence of the observational estimators. Then second-order effects are negligible (and Assumption 2 holds asymptotically) provided that $b_{n}^{2}/\sqrt{n}\rightarrow 0$. This implies that Assumption 2 holds when the asymptotic bias $b_{n}$ is proportional to, smaller or even much larger than the asymptotic variance, provided that $b_{n}$ does not grow at rate faster than $n^{1/4}$.
@@
Gechter et al. (2024) study where to run an experiment. For illustration, consider two sites $j\in\{1,2\}$ with site-specific ATEs $\theta_{j}$ and target the cross-site average
$\tau(\theta)=\tfrac{1}{2}(\theta_{1}+\theta_{2})=\omega^{\top}\theta,\qquad\omega=\tfrac{1}{2}(1,1)^{\top}.$
Let $\tilde{\theta}^{\mathrm{obs}}=(\tilde{\theta}^{\mathrm{obs}}_{1},\tilde{\theta}^{\mathrm{obs}}_{2})^{\top}$ denote observational estimates obtained in Gechter et al. (2024) from a structural model and potentially biased due to misspecification: $\mathbb{E}[\tilde{\theta}^{\mathrm{obs}}]-\theta=(b_{1},b_{2})^{\top}.$ A budget constraint allows an experiment in only one site, so that $\mathcal{E}\in\{\{1\},\{2\}\}$. If site $j$ is chosen, we obtain an unbiased $\tilde{\theta}^{\mathrm{exp}}_{j}$; the other site $k\neq j$ remains observational. Given $\mathcal{E}$,
$\widehat{\tau}_{\gamma}=\tfrac{1}{2}\!\left[\gamma_{j}\,\tilde{\theta}^{\mathrm{exp}}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j}+\tilde{\theta}^{\mathrm{obs}}_{k}\right],\qquad k\neq j.$
@@
Example 3 (Supply or demand experiment with non linear target).
Bergquist and Dinerstein (2020) conduct demand and supply experiments in food markets in Kenya. Suppose here we are interested in similar applications in Uganda. For exposition, consider a basic linear demand and supply
@@
with $\beta_{D}\neq-\beta_{S}$. Several estimands may be of interest; one of such estimands is the effect of a tariff $t$ on prices
$\tau\equiv\frac{\beta_{S}}{\beta_{D}+\beta_{S}}\,t.$
Let $(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top}$ denote baseline estimates from Kenya, that may lack external validity.
Due to cost constraints, $\mathcal{E}\in\{\{D\},\{S\}\},$ i.e., we can learn either demand (estimating $\tilde{\beta}_{D}^{\mathrm{exp}}$, with randomized price discounts) or supply (estimating $\tilde{\beta}_{S}^{\mathrm{exp}}$, by introducing regulatory cost shocks on firms). Let
@@
with $\theta$ and $\tilde{\theta}$ denoting the parameter and estimator here rescaled by the square-root of the sample size, and $b$ capturing bias. Our goal is to choose whether to conduct a supply or demand experiment in Uganda to estimate $\tau$. For given estimators $\hat{\theta}$ that combine the estimates from Kenya with the chosen experimental estimate from Uganda under local asymptotics described in Section 5.1
$\tau-\widehat{\tau}=\underbrace{\frac{1}{\sqrt{n}}\check{\omega}(\mathrm{plim}(\tilde{\beta}^{\mathrm{obs}}))^{\top}(\theta-\widehat{\theta})}_{\text{main first order effect}}+\underbrace{o_{p}\left(\frac{1}{n^{1/2}}\right)}_{\text{small higher order effects}},\qquad\check{\omega}(\beta)=\frac{t}{(\beta_{D}+\beta_{S})^{2}}\begin{pmatrix}-\beta_{S}\\ \beta_{D}\end{pmatrix}$,
where we can replace $\check{\omega}$ with its consistent counterpart $\check{\omega}(\tilde{\beta}^{\mathrm{obs}})$. Our goal is to choose between a supply or demand experiment accounting for first-order bias (and variance) effect.
@@ 3 Robust experimental design
Next, we introduce an experimental design focusing on the mean-squared error (MSE) of the estimator $\hat{\tau}$, defined, for given bias level $b$ and design $(\mathcal{E},\Sigma(\mathcal{E}))$ as
$\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)\;=\;\mathbb{E}_{\mathcal{E},\Sigma,b}\!\left[(\widehat{\tau}_{\gamma}-\tau)^{2}\right],$
where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the data-generating process implied by $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$. Section 5.2 shows that our results for MSE also extend when minimizing confidence intervals’ length.
@@
$\mathcal{B}(B)\equiv\{b:\ \|b\|_{\infty}\leq B\},$
with $B\geq 0$ an upper bound on the largest coordinate-wise bias. In our framework, a key challenge is that biases may arise from multiple parameters and uncertainty sets defined by different norms may be possible. A key insight here is that the $\ell_{\infty}$ choice has a desirable property relative to $\ell_{1}$ or $\ell_{2}$: it does not force a trade-off across coordinates (a large bias in one component need not be “offset” by a small bias elsewhere), which is attractive when biases may be positively correlated across parameters. The drawback is that worst-case solutions depend on the radius $B$ which may be unknown in practice.
@@
$\mathcal{R}(\mathcal{E},\Sigma,\gamma)\;\equiv\;\sup_{B\geq 0}\ \frac{\displaystyle\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}{\displaystyle\mathrm{MSE}^{*}(B)},$
defined adaptation regret by Armstrong et al. (2024) (building in turn on Tsybakov, 1998).
There are two key distinctions relative to such references using the adaptation regret. First (and importantly) we optimize over the design itself rather than fixing it ex ante; consequently, regret is defined relative to an oracle that jointly chooses both the estimator and the design. Second, we allow biases to arise across multiple parameters (rather than a single one), which necessitates a careful specification of the uncertainty set; we adopt an $\ell_{\infty}$-ball, which, as we show yields transparent and computationally-attractive solutions.
@@ 3.1 Optimal design
A natural question is whether the adaptation regret provides natural trade-offs between the variance and the bias. To do so, denote $\mathbb{V}_{\mathcal{E},\Sigma}(\cdot)$ the variance operator for given $(\mathcal{E},\Sigma)$.
@@
$\alpha(\mathcal{E},\Sigma,\gamma)\;\equiv\;\mathbb{V}_{\mathcal{E},\Sigma}(\hat{\tau}_{\gamma}),\qquad\alpha^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\ \min_{\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})}\ \min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\alpha(\mathcal{E},\Sigma,\gamma),$ (2)
@@
The variance regret denotes the ratio between the variance of a given design and estimator relative to the smallest achievable variance. We provide a similar definition for the bias. Let $|1-\gamma|$ define the absolute value of each entry of the vector $1-\gamma$.
@@
$\qquad\beta(\mathcal{E},\gamma)\;\equiv\;\Big(\,\|\omega\|_{1}-\|\omega_{\,\mathcal{E}}\|_{1}+\,|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\,\Big)^{\!2},\quad\beta^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\Big(||\omega||_{1}-\|\omega_{\mathcal{E}}\|_{1}\Big)^{2},$
@@
[…]
An ideal design, would set the variance equal to $\alpha^{\star}$ and bias equal to $\beta^{\star}$. Unfortunately, this may be infeasible as it might require different choices of experiments and estimators to achieve one or the other. This raises the question of how to trade-off these two components.
This trade-off is formally characterized in our main theorem below.
@@
Theorem 1 provides a simple expression of the regret which does not require researchers to specify $B$. The key insight is that the adaptation regret in our context, leads to a (strictly) quasi-convex objective function, whose worst-case solution is the maximum between the variance and the bias regret. This provides a simple and intuitive trade-off between the two.
Specifically, provided that the class of designs $\mathcal{D}$ is sufficiently flexible (outside boundary solutions), it follows that at the optimum we equalize the variance and bias regret as in Figure 1
@@
In practice, this amount to allocate sample size to noisier treatment arms when the variance dominates, and select treatment arms with largest sensitivity when the bias dominates.
@@
Theorem 1 provides us an immediate solution to the optimal design problem. We compute the estimator, the optimal variance and the experimental choice via backward induction:
$\gamma^{\star}(\mathcal{E},\Sigma)$ $\in\mathrm{arg}\min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma,\gamma)}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma)}{\beta^{\star}}\right\}$
$\Sigma^{\star}(\mathcal{E})$ $\in\mathrm{arg}\min_{\Sigma\in\mathcal{G}(\mathcal{E})}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma,\gamma^{\star}(\mathcal{E},\Sigma))}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma^{\star}(\mathcal{E},\Sigma))}{\beta^{\star}}\right\}$
$\mathcal{E}^{\star}$ $\in\mathrm{arg}\min_{\mathcal{E}\in\mathcal{S}}\,\max\!\left\{\frac{\alpha(\mathcal{E},\Sigma^{\star}(\mathcal{E}),\gamma^{\star}(\mathcal{E},\Sigma^{\star}(\mathcal{E})))}{\alpha^{\star}},\ \ \frac{\beta(\mathcal{E},\gamma^{\star}(\mathcal{E},\Sigma^{\star}(\mathcal{E})))}{\beta^{\star}}\right\}.$
That is, we first choose $\gamma^{\star}$ for each design $\mathcal{E},\Sigma(\mathcal{E})$. We then choose $\Sigma$, whose choice depends on $\beta$ through the shrinkage weights. We finally choose the set of experiments $\mathcal{E}$.
3.2 Intuition with two parameters
To build further intuition, it is useful to study the two-parameter model.
Setting 2 (Illustration with two-parameter model).
Let $\theta=(\theta_{1},\theta_{2})^{\top}$, and mutually independent observational and experimental estimates $\tilde{\theta}^{\mathrm{obs}}_{j}-\theta_{j}\sim\mathcal{N}(b_{j},\ \sigma_{j}^{2}),\quad\tilde{\theta}^{\exp}_{j}-\theta_{j}\sim\mathcal{N}(0,\ v_{j}^{2}),\quad j=1,2.$ For simplicity $\sigma_{j}^{2},v_{j}^{2}$ are fixed. We may run one experiment due to budget constraints: $\mathcal{E}\in\{\{1\},\{2\}\}$. If we pick $j$, we estimate $\hat{\theta}_{j}=\gamma_{j}\,\tilde{\theta}^{\exp}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j},\quad\hat{\theta}_{-j}=\tilde{\theta}^{\mathrm{obs}}_{-j},$ and, to first order, $\tau(\theta)-\tau(\hat{\theta})=\omega_{1}(\theta_{1}-\hat{\theta}_{1})+\omega_{2}(\theta_{2}-\hat{\theta}_{2}).$
How do the variance and bias regret look like in this easier scenario? We can write
@@ 3.2 Intuition with two parameters
The very first step is optimizing over the estimator (assuming here for simplicity $\gamma$ can take any value). The following characterizes the regret optimal $\gamma^{\star}$ (see Appendix C.1).

### v3

@@ 2 Baseline scenario: problem description
We assume that researchers have access to estimators (and their variance) of $\theta$ denoted as $\tilde{\theta}^{\mathrm{obs}}\in\mathbb{R}^{p}$. We impose no restrictions on how $\tilde{\theta}^{\mathrm{obs}}$ is formed: it can based on arbitrary exclusion restrictions implied by an economic or statistical model. However, $\tilde{\theta}^{\mathrm{obs}}$, defined as observational estimates, have biases collected in a vector $b\in\mathbb{R}^{p}$ unknown at the experimental design stage (see Remark 2). Examples include an estimate from an instrumental variable regression that may fail the exclusion restriction, or an experimental estimate from a country different from the one of interest that may lack external validity. Researchers can then collect experimental evidence for a subset of parameters, producing estimates which are unbiased by design.
@@
Here, $(\mathcal{E},\Sigma(\mathcal{E}))$ characterizes the experimental design, encoding which parameters are learned and with what precision. When clear, we omit the subscript $\Sigma$ on $\tilde{\theta}^{\mathrm{exp}}$. We assume $\Sigma(\mathcal{E})$ is known, which is standard in experimental planning (Gerber and Green, 2012); when $\Sigma$ is unknown, we can interpret $\Sigma$ as the expected variance-covariance matrix. [^7] In this case, we interpret the MSE in Equation (2) as the expected MSE under a common expectation over the experimental variance-covariance matrix. It is also possible to extend our framework by allowing $\Sigma$ to be a function of $\tilde{\theta}^{\mathrm{obs}}$ provided that $\Sigma(\tilde{\theta}^{\mathrm{obs}})\rightarrow_{p}\Sigma(\mathrm{plim}(\tilde{\theta}^{\mathrm{obs}}))$, with $\Sigma(\mathrm{plim}(\tilde{\theta}^{\mathrm{obs}}))$ denoting the common expectation of the variance-covariance experimental matrix. In this case our analysis should be interpreted in an asymptotic regime similar to what discussed in Section 5.3.
@@
We write $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$ indicating $\mathcal{D}$ the set of feasible designs (i.e., choice of $\mathcal{E}$ and $\Sigma(\mathcal{E})$) and estimators (i.e., $\gamma$). The set $\mathcal{D}$ can be arbitrary as long as (i) each feasible $\Sigma$ has uniformly bounded entries and is strictly positive definite; (ii) $\gamma_{j}=1$ for all $j\in\mathcal{E}$ is one admissible choice (among potentially others).
We let $(\mathcal{E},\Sigma(\mathcal{E}))$ be within arbitrary constraint sets that may encode feasibility or budget constraints. In our framework, restrictions on the experiments researchers can run $\mathcal{E}$ and on their precision $\Sigma(\mathcal{E})$ can take any desired form. In most applications, we think of such constraints as arising from fixed costs of running an experiment and/or constraints on their power (Duflo et al., 2007; Athey and Imbens, 2017; List et al., 2011). We assume that $\Sigma(\mathcal{E})$ is uniformly bounded, which implies that once we commit to learn a set of experimental parameters $\tilde{\theta}_{\mathcal{E}}^{\mathrm{exp}}$, their variance is finite, and strictly positive definite, therefore assuming that the variance is bounded away from zero. [^8] For this latter condition to hold in an asymptotic framework with growing sample size, $\theta$ and $\tilde{\theta}$ can be defined without loss as parameters of interest after appropriately rescaling by the square-root of the sample size; in this case $\Sigma$ denotes the asymptotic variance. See Section 5.3 for details. The shrinkage estimator can be unrestricted (i.e., $\gamma$ can take any value on the real line) or restricted (when for example researchers aim to only use experimental evidence when available, setting $\gamma_{j}=1$ for $j\in\mathcal{E}$).
Example 1 (Structure of $\Sigma$).
Because observational and experimental moments are typically computed from disjoint samples, in practice, it is natural to model the covariance as block diagonal, $\Sigma=\operatorname{diag}(\Sigma_{\text{obs}},\Sigma_{\text{exp}})$. The first block corresponds to observational estimates (fixed and not subject to design choices), while the second corresponds to experimental estimates, and is a function of the sample size allocated to treatment arms (that we may optimize). This is a special case of our framework which can be encoded in the constraint set $\mathcal{D}$. ∎
Remark 1 (Some unbiased observational estimates).
In some applications, researchers may assume that some of the observational study estimates are not biased ($b_{j}=0$ for some $j$). This is possible to accommodate as a special case of our general framework in Section 4. ∎
Main design problem
@@
Choosing $\gamma$ (step 1) for a fixed design follows a long-standing tradition in econometrics and statistics (Armstrong et al., 2024; Andrews and Shapiro, 2021; Donoho, 1994; Athey et al., 2025). We study this question here in combination with the experimental design problem, step 2 and 3, which, as we show will change the decision problem. The challenge is that performance depends on the bias $b$, unknown when designing the experiment.
Properties of $\tau$
A key feature of our framework is that researchers explicitly parameterize $\tau(\cdot)$, thereby pre-specifying and making transparent which biases drive the design problem; this feature clarifies the role of the experiment in complementing existing evidence. The sensitivity of $\tau$ to $\theta$, captured by its gradient $\omega(\theta)=\partial\tau(\theta)/\partial\theta$, determines how (to first order) the bias propagates onto the final estimand. For simplicity, we impose exact linearity, $\tau(\theta)=\omega^{\top}\theta$; all results extend to smooth, nonlinear $\tau$ via a first-order expansion. When $\tau$ is unknown, Section 5.1 studies settings where researchers only partially identify $\tau$.
@@
[…]
@@
Gechter et al. (2024) study where to run an experiment. For illustration, consider two sites $j\in\{1,2\}$ with site-specific ATEs $\theta_{j}$ and target the cross-site average $\tau(\theta)=\omega_{1}\theta_{1}+\omega_{2}\theta_{2},$ with $\omega$ denoting the population share in each site. Let $\tilde{\theta}^{\mathrm{obs}}=(\tilde{\theta}^{\mathrm{obs}}_{1},\tilde{\theta}^{\mathrm{obs}}_{2})^{\top}$ denote observational estimates obtained in Gechter et al. (2024) from a structural model and potentially biased due to misspecification: $\mathbb{E}[\tilde{\theta}^{\mathrm{obs}}]-\theta=(b_{1},b_{2})^{\top}.$ A budget constraint allows an experiment in only one site, so that $\mathcal{E}\in\{\{1\},\{2\}\}$. If site $j$ is chosen, we obtain an unbiased $\tilde{\theta}^{\mathrm{exp}}_{j}$; the other site $k\neq j$ remains observational. Given $\mathcal{E}$,
$\widehat{\tau}_{\gamma}=\omega_{j}\gamma_{j}\,\tilde{\theta}^{\mathrm{exp}}_{j}+\omega_{j}(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j}+\omega_{k}\tilde{\theta}^{\mathrm{obs}}_{k},\qquad k\neq j.$
@@
Example 4 (Experiment with non linear target and equilibrium model).
Bergquist and Dinerstein (2020) conduct demand and supply experiments in food markets in Kenya. Suppose here we are interested in similar applications in Uganda. For exposition, consider a stylized linear demand and supply model (similar reasoning applies to more complex models)
@@
with $\beta_{D}\neq-\beta_{S}$. Several estimands may be of interest; one of such estimands is the effect of a tariff $t$ on prices $\tau\equiv\frac{\beta_{S}}{\beta_{D}+\beta_{S}}\,t.$ Let $(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top}$ denote baseline estimates from Kenya, that may lack external validity.
Due to fixed costs of each experiment, $\mathcal{E}\in\{\{D\},\{S\}\},$ i.e., we can learn either demand (estimating $\tilde{\beta}_{D}^{\mathrm{exp}}$, with randomized price discounts) or supply (estimating $\tilde{\beta}_{S}^{\mathrm{exp}}$, by introducing regulatory cost shocks on firms). Let
@@
with $\theta$ and $\tilde{\theta}$ denoting the parameter and estimator here rescaled by the square-root of the sample size, and $b$ capturing bias. Our goal is to choose whether to conduct a supply or demand experiment in Uganda to estimate $\tau$. For given estimators $\hat{\theta}$ that combine the estimates from Kenya with the chosen experimental estimate from Uganda under local asymptotics described in Section 5.3 and Remark 6
$\tau-\widehat{\tau}=\underbrace{\frac{1}{\sqrt{n}}\omega^{\top}(\theta-\widehat{\theta})}_{\text{main first order effect}}+\underbrace{o_{p}\left(\frac{1}{n^{1/2}}\right)}_{\text{small higher order effects}},\hskip 18.49988pt\omega=\frac{t}{(\mathrm{plim}(\tilde{\beta}_{D}^{\mathrm{obs}}+\tilde{\beta}_{S}^{\mathrm{obs}}))^{2}}\begin{pmatrix}-\mathrm{plim}(\tilde{\beta}_{S}^{\mathrm{obs}})\\ \mathrm{plim}(\tilde{\beta}_{D}^{\mathrm{obs}})\end{pmatrix}$,
where we can replace $\mathrm{plim}(\tilde{\beta}^{\mathrm{obs}})$ with its consistent counterpart $\tilde{\beta}^{\mathrm{obs}}$.
@@ 3 Robust experimental design
Next, we introduce an experimental design focusing on the mean-squared error (MSE) of the estimator $\hat{\tau}$. We consider confidence interval length in Section 5.1. The MSE is defined, for given bias level $b$, design $(\mathcal{E},\Sigma(\mathcal{E}))$, and choice of weights $\gamma$ as
$\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)\;=\;\mathbb{E}_{\mathcal{E},\Sigma,b}\!\left[(\widehat{\tau}_{\gamma}-\tau)^{2}\right],$ (2)
where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the design $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$.
@@
$\mathcal{B}(B)\equiv\{b:\ \|b\|_{\infty}\leq B\},$ (3)
with $B\geq 0$ an upper bound on the largest coordinate-wise bias. In our framework, a key challenge is that biases may arise from multiple parameters. Here, for exposition we first focus on the $\ell_{\infty}$-ball given its interpretability, while Section 4 generalizes the framework to arbitrary norms. The $\ell_{\infty}$ norm imposes symmetric restrictions on the biases and it does not force a trade-off across coordinates (a large bias in one component need not be “offset” by a small bias elsewhere), which is attractive when biases may be positively correlated across parameters. The drawback is that worst-case solutions depend on the radius $B$ which may be unknown in practice.
@@
$\mathcal{R}(\mathcal{E},\Sigma,\gamma)\;\equiv\;\sup_{B\geq 0}\ \frac{\displaystyle\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(\mathcal{E},\Sigma,\gamma)}{\displaystyle\mathrm{MSE}^{*}(B)},$ (4)
defined adaptation regret by Armstrong et al. (2024) and Tsybakov (1998) for estimation under fixed designs (see Remark 3 for a comparison with our experimental design problem).
Remark 2 (Adaptation to $B$).
We consider worst-case adaptation regret over $B$ because researchers typically do not know $B$ before running an experiment and learning $B$ would require running pilots for each design in the menu of feasible experiments. Also, Armstrong and Kolesár (2018) show that pointwise adaptation to unknown misspecification levels is impossible in general, motivating the adaptation regret. Note however that if researchers know that the misspecification level $B$ lies in a bounded interval $B\in[0,\bar{B}]$, our characterization extends by simply restricting the worst-case over $B$ to this interval; see Section 5.2.∎
@@ 3.1 Optimal design
Our main results show that the adaptation regret provides natural trade-offs between the variance and the bias. To do so, denote $\mathbb{V}_{\mathcal{E},\Sigma}(\cdot)$ the variance operator for given $(\mathcal{E},\Sigma)$.
@@
$\alpha(\mathcal{E},\Sigma,\gamma)\;\equiv\;\mathbb{V}_{\mathcal{E},\Sigma}(\hat{\tau}_{\gamma}),\qquad\alpha^{\star}\;\equiv\;\min_{(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}}\alpha(\mathcal{E},\Sigma,\gamma),$ (5)
@@
The variance regret denotes the ratio between the variance of a given design and estimator relative to the smallest achievable variance. We provide a similar definition for the bias.
@@
$\qquad\beta(\mathcal{E},\gamma)\;\equiv\;\Big(\,\|\omega\|_{1}-\|\omega_{\,\mathcal{E}}\|_{1}+\,|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\,\Big)^{\!2},\quad\beta^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\Big(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1}\Big)^{2},$
@@
To gain insight, note that $B^{2}\beta=B^{2}\Big(\|\omega\|_{1}-\|\omega_{\mathcal{E}}\|_{1}\Big)^{2}$ denotes the worst-case bias for a given design, value $B^{2}$ and shrinkage estimator. $B^{2}\beta^{\star}$ denotes its minimizer.
An ideal design would set the variance equal to $\alpha^{\star}$ and bias equal to $\beta^{\star}$. Unfortunately, this may be infeasible as it might require different choices of experiments and estimators to achieve one or the other. This raises the question of how to trade-off these two components.
This trade-off is formally characterized in our main theorem below. We will use the convention throughout that 0/0 = 1.
@@
Theorem 1 offers a simple and novel expression of the regret which does not require researchers to specify $B$. As we show below, Theorem 1 will significantly simplify the optimization program. The key insight is that the adaptation regret for linear estimators and arbitrary designs leads to a (strictly) quasi-convex objective function, whose worst-case solution is the maximum between the variance and the bias regret. This provides a simple and intuitive trade-off between the two. Specifically, provided that the class of designs $\mathcal{D}$ is sufficiently flexible (outside boundary solutions), it follows that at the optimum we equalize
@@
That is, we allocate sample size to noisier treatment arms when the variance dominates, and select treatment arms with largest sensitivity when the bias dominates.
@@
Remark 3 (Comparison with previous literature on adaptation regret).
Previous literature has studied adaptation regret to combine an unbiased and a biased estimator for a single parameter $\theta\in\mathbb{R}$ under a fixed design (Armstrong et al., 2024; Tsybakov, 1998). There are three key distinctions from this literature. First (and foremost), we optimize over the design itself rather than taking it as given; regret is defined relative to an oracle that chooses both the estimator and the design ex ante. Second, we allow biases to arise across multiple parameters (rather than a single one), and do not require that researchers have access to an unbiased estimator for each component of $\theta$. Third, we study a linear class of estimators, instead of arbitrary nonlinear rules. [^10] In the univariate bounded-normal-mean problem, Donoho et al. (1990) show that linear estimators lose no more than 25% efficiency relative to the optimal nonlinear estimator.
As a consequence, the characterization in Theorem 1 differs from Armstrong et al. (2024), where regret is not quasi-convex in the bias radius and the optimal estimator solves a minimax problem via a least-favourable prior; see Appendix D. Their minimax solution for combining two estimators is not applicable here once we optimize over the choice of experiments, sample allocation and multiple estimators, since searching over nonlinear shrinkage rules becomes computationally demanding. By focusing on linear estimators, we obtain a tractable solution that also extends to common method-of-moments, see Section 4. ∎
Example 5 (Intuition with two parameters).
To build further intuition, it is useful to study the two-parameter model. Let $\theta=(\theta_{1},\theta_{2})^{\top}$, and mutually independent observational and experimental estimates $\tilde{\theta}^{\mathrm{obs}}_{j}-\theta_{j}\sim\mathcal{N}(b_{j},\ \sigma_{j}^{2}),\quad\tilde{\theta}^{\exp}_{j}-\theta_{j}\sim\mathcal{N}(0,\ v_{j}^{2}),\quad j=1,2.$ For simplicity $\sigma_{j}^{2},v_{j}^{2}$ are fixed. We may run one experiment due to budget constraints: $\mathcal{E}\in\{\{1\},\{2\}\}$. If we pick $j$, we estimate $\hat{\theta}_{j}=\gamma_{j}\,\tilde{\theta}^{\exp}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j},\quad\hat{\theta}_{-j}=\tilde{\theta}^{\mathrm{obs}}_{-j},$ and, to first order, $\tau(\theta)-\tau(\hat{\theta})=\omega_{1}(\theta_{1}-\hat{\theta}_{1})+\omega_{2}(\theta_{2}-\hat{\theta}_{2}).$ We can write
@@
The very first step is optimizing over the estimator (assuming here for simplicity $\gamma$ can take any value). The following characterizes the regret optimal $\gamma^{\star}$ (see Appendix B.1).